Randall and Amyra left school at the same time and began walking in opposite directions. Randall walked at a rate of 3.6 mph and amyra walked at a rate of 4.2 mph. How far apart were the after 10 minutes
Answer by Paul(988)

Since its miles per hour
Divide 10 minutes by 60 minutes=1/6
3.6x+4.2x=1/6
simplfy
x=.03
Hence they will be 0.03miles apart after 10 minutes.
Paul.